OpenStudy (anonymous):

Use the Factor Theorem to determine whether the first polynomial is a factor of the second polynomial. x - 5; 3x^2 + 7x + 40

OpenStudy (anonymous):

YES right?

OpenStudy (anonymous):

x-5 is not a factor

OpenStudy (anonymous):

so no?

OpenStudy (anonymous):

nope

OpenStudy (zehanz):

x-5 is a factor if substituting 5 in the second gives 0: 3*5² + 7*5 + 40 = 75 + 35 + 40 = 150 <> 0, so NO.

OpenStudy (raden):

this cant be factored check the value of D (discriminant) = b^2-4ac if D < 0, it has complex roots